OpenStudy (anonymous):

4^2x = 32^1/2 Solve for x.

OpenStudy (campbell_st):

same as before... use a base of 2 \[(2^2)^{2x} = (2^5)^{\frac{1}{2}}\]

OpenStudy (anonymous):

\[4^{2x}=32^{1/2}\] \[32^{1/2} = \sqrt{32} = 4\sqrt{2}\] \[2x = \frac{ \log(4\sqrt{2)} }{ \log(4) } \] just divide by 2 now.

OpenStudy (campbell_st):

use the index laws... to simplify the powers and now they are the same base... just equate the powers

OpenStudy (anonymous):

Thanks so much you guys! That was much simpler than how my teacher explained it. :)